An office supply company sells Q permanent markers per year at $P per marker. The price-demand equation for these markers is P=10-0.001Q. The total cost of manufacturing Q permanent markers is TC(Q)=5000+2Q.


A. What is the company's maximum profit?B. What should the company charge for the markers, and how many markers should be produce to maximize the profit?

1
Expert's answer
2021-07-19T05:48:30-0400

A.

"Profit=Revenue-Cost"

"P(Q)=R(Q)-TC(Q)"


"R(Q)=p(Q)Q=(10-0.001Q)Q"

"=10Q-0.001Q^2"

"P(Q)=10Q-0.001Q^2-(5000+2Q)"

"=8Q-0.001Q^2-5000"


"P(Q)=8Q-0.001Q^2-5000"

B.


"Q\geq0, p(Q)\geq0"

"10Q-0.001Q^2\geq0"

"0\leq Q\leq 10000"

"P'(Q)=(8Q-0.001Q^2-5000)'"

"=8-0.002Q"

Find critical number(s)


"P'(Q)=0=>8-0.002Q=0=>Q=4000"

If "0\leq Q\leq4000, P'(Q)>0, P(Q)" increases.


If "4000\leq Q\leq10000, P'(Q)<0, P(Q)" decreases.

The function "P(Q)" has a local maximum at "Q=4000."


Since the function "P(Q)" has the only extremum on "[0, 10000]," then the function "P(Q)" has the absolute maximum at "Q=4000."


"P(4000)=8(4000)-0.001(4000)^2-5000"

"P(4000)=\$11000"

"p(4000)=10-0.001(4000)"

"p(4000)=\$6"